OpenStudy (anonymous):

12-4x≤20 Which of these shows the solution to this inequality? A B C D

OpenStudy (anonymous):

\[12 - 4x \le 20\] Subtract 12 from 12 and 20 \[12 - 4x \le 20\] -12 -12 _______________ \[-4x\le 20\] (Divide each side by -4) \[-4x\le 20\] ____ ____ -4 -4 \[x\le\] -5 (When dividing by negatives, the inequality sign flips) \[x\ge-5\]